# Leveraging Second-Order Curvature for Efficient Learned Image Compression: Theory and Empirical Evidence

### v1

@@ Abstract
Training learned image compression (LIC) models entails navigating a challenging optimization landscape defined by the fundamental trade-off between rate and distortion. Standard first-order optimizers, such as SGD and Adam, struggle with gradient conflicts arising from competing objectives, leading to slow convergence and suboptimal rate–distortion performance. In this work, we demonstrate that a simple utilization of a second-order quasi-Newton optimizer, SOAP, dramatically improves both training efficiency and final performance across diverse LICs. Our theoretical and empirical analyses reveal that Newton preconditioning inherently resolves the intra-step and inter-step update conflicts intrinsic to the R–D objective, facilitating faster, more stable convergence. Beyond acceleration, we uncover a critical deployability benefit: second-order trained models exhibit significantly fewer activation and latent outliers. This substantially enhances robustness to post-training quantization. Together, these results establish second-order optimization, achievable as a seamless drop-in replacement of the imported optimizer, as a powerful, practical tool for advancing the efficiency and real-world readiness of LICs.
@@ I Introduction
  • We empirically demonstrate that the second-order optimizer substantially reduces both training steps and wall-clock time required to achieve the same performance, while simultaneously improving the final rate–distortion performance compared to first-order optimizers when trained for the same number of steps. (Sec. IV)
  • We provide theoretical and empirical evidence that the second-order optimizer aligns gradients update from competing R–D loss terms and from consecutive update steps, enabling more effective optimization trajectories. (Sec. V)
@@ II-A Advances in Learned Image Compression Architectures
Learned image compression methods are typically developed within the nonlinear transform coding framework [4, 30], aiming to balance the bit rate (R) and the distortion (D). A wide variety of architectures have been explored to enhance the expressive power of transforms. Examples include residual networks [38, 13], deformable convolutions [24], designs based on frequency decomposition [26, 61], invertible neural networks [98, 12, 28], and contextual clustering [78]. Recently, transformers and Mamba architectures have gained traction, offering strong performance gains [57, 113, 114, 47, 79, 50, 80, 97, 105, 21, 58]. On the probabilistic modeling side, research has sought to design more accurate entropy models for the latent space. This includes hierarchical priors [ballé2018variational, 40, 17, 54, 111], autoregressive models operating over spatial [65] or channel dimensions [67], as well as hybrid models that capture joint spatial–channel dependencies [44, 59]. Further refinements make use of checkerboard-based decoding [39, 24], codebooks [112], distribution [25, 106], and (lattice) vector quantization [108, 22, 48, 99].
Another body of work seeks to reduce training and inference cost while maintaining compression quality. Notable examples include slimmable sub-networks [89], variable-bit-rate codecs [33, 45, 77], and knowledge distillation strategies [24]. Lightweight decoding has been pursued through shallow or linear decoders [102], while new loss formulations such as causal context [35] or latent decorrelation penalties [1] have also been proposed. Recent studies further introduce rate–distortion–complexity analysis [66, 29, 32, 36], explicitly incorporating computational cost into the optimization objective.
@@ II-B Training Dynamics Approaches for LIC
Although rate–distortion training is often cast as minimizing a scalarized loss $R+\lambda D$, it is fundamentally a multi-objective problem: decreasing rate typically worsens distortion, and vice versa. This observation motivates the adoption of multi-objective optimization (MOO) techniques, which are designed to handle multiple conflicting criteria. One influential MOO approach is the Multiple Gradient Descent Algorithm (MGDA) [15, 86]. MGDA determines an update direction by combining gradients from different objectives with non-negative weights that minimize the squared norm of their sum, subject to a simplex constraint. The resulting direction guarantees improvement for all objectives simultaneously. MOO methods have seen wide adoption in multi-task learning [56, 95, 41], where they are used to balance competing gradients across tasks and mitigate conflicts during training.
@@
Our Perspective. Our work aligns with this emerging line of training-dynamics-based methods and is closely related to Balanced-RD, CMD-LIC, and AuxT. Balanced-RD explicitly regulates the interaction between rate and distortion gradients to promote stable convergence in the R–D trade-off. CMD-LIC accelerates optimization by reducing training space dimensions. AuxT introduces constraints on energy compaction and feature decorrelation to improve convergence behavior.
@@ V-D Empirical Validation
These empirical findings substantiate our central claim: SOAP’s performance gains arise from resolving both intra- and inter-step gradient conflicts in rate–distortion optimization.
@@ VI-C Empirical Validation
These empirical findings support the theoretical perspective in Sec. VI: By coupling channels via Newton-like scaling and rotations, SOAP redistributes variance across directions rather than concentrating it in a few, preventing outlier formation. This yields more regular feature statistics, improving entropy modeling and stabilizing activations, with the downstream benefit of enhanced PTQ robustness. Thus, SOAP not only accelerates training and improves R–D trade-offs but also produces models that are substantially easier to deploy on constrained hardware.
@@ A-D Limitations of Adam for Gradient Alignment
Adam is powerful and widely used, but its effectiveness is limited by its diagonal preconditioner. Because it scales coordinates independently, it cannot exploit off–diagonal curvature that encodes interactions among parameters—precisely what is needed to resolve non–axis-aligned gradient conflicts in multi-objective settings such as rate–distortion (R–D) optimization. The following proposition formalizes a standard local approximation behind this limitation, following [69, 62].
@@ A-E Adam’s Gradient Conflict in a Simplified R-D Setting
Note (positive per–coordinate scalings). The orthogonality conclusion in Proposition 3 is unchanged if Adam’s elementwise normalization introduces arbitrary positive scalings that are independent of the signs of the initialized weights. Concretely, let
@@ A-F Adam’s Behaviour Near a Nondegenerate Optimum
    Frozen second moments. Adam’s second–moment accumulator and damping are (locally) stationary, yielding a fixed positive diagonal matrix $D\succ 0$ and a scalar $c>0$ (absorbing stepsize, bias correction, $\varepsilon$). Hence the Adam update is the diagonally preconditioned step [46, 81, 104]
@@
Since $B$ is symmetric positive definite, inter-step cosines in the qq–space admit closed forms; cosines in the original coordinates are equivalent up to constants depending only on $\kappa(D)$ (the condition number of D) [76].
@@ A-F1 Inter-step cosine for Adam
the exact inter-step cosine in the qq–space is
@@
  • Oscillation threshold. If $\eta\,\mu_{1}(u_{t})>1$, the numerator in equation 58 is negative and $\mathcal{S}_{q}<0$. Hence whenever $\eta\,\lambda_{\max}(B)>1$, there exist directions with negative inter-step cosine (flip–flop behaviour).
@@ A-F2 Intra-step cosine for Adam near the optimum
Remarks Nonzero momentum ($\beta_{1}>0$) produces a linear two-term recurrence in the qq–space; all qualitative conclusions above persist with $B$ replaced by an $O(1)$ affine function of $B$. If $D$ evolves slowly rather than remaining fixed, equation 58–equation 60 apply between preconditioner refreshes with the current $D_{t}$.

### v2

@@ Abstract
Training learned image compression (LIC) models entails navigating a challenging optimization landscape defined by the fundamental trade-off between rate and distortion. Standard first-order optimizers, such as SGD and Adam, struggle with gradient conflicts arising from competing objectives, leading to slow convergence and suboptimal rate-distortion performance. In this work, we demonstrate that a simple utilization of a second-order quasi-Newton optimizer, SOAP, dramatically improves both training efficiency and final performance across diverse LICs. Our theoretical and empirical analyses reveal that Newton preconditioning inherently resolves the intra-step and inter-step update conflicts intrinsic to the R-D objective, facilitating faster, more stable convergence. Beyond acceleration, we uncover a critical deployability benefit: second-order trained models exhibit significantly fewer activation and latent outliers. This substantially enhances robustness to post-training quantization. Together, these results establish second-order optimization, achievable as a seamless drop-in replacement of the imported optimizer, as a powerful, practical tool for advancing the efficiency and real-world readiness of LICs.
@@ I Introduction
  • We empirically demonstrate that the second-order optimizer substantially reduces both training steps and wall-clock time required to achieve the same performance, while simultaneously improving the final rate-distortion performance compared to first-order optimizers when trained for the same number of steps. (Sec. IV)
  • We provide theoretical and empirical evidence that the second-order optimizer aligns gradients update from competing R-D loss terms and from consecutive update steps, enabling more effective optimization trajectories. (Sec. V)
@@ II-A Advances in Learned Image Compression Architectures
Learned image compression methods are typically developed within the nonlinear transform coding framework [4, 30], aiming to balance the bit rate (R) and the distortion (D). A wide variety of architectures have been explored to enhance the expressive power of transforms. Examples include residual networks [38, 13], deformable convolutions [24], designs based on frequency decomposition [26, 61], invertible neural networks [97, 12, 28], and contextual clustering [78]. Recently, transformers and Mamba architectures have gained traction, offering strong performance gains [57, 113, 114, 47, 79, 50, 96, 104, 21, 58]. On the probabilistic modeling side, research has sought to design more accurate entropy models for the latent space. This includes hierarchical priors [ballé2018variational, 40, 17, 54, 111], autoregressive models operating over spatial [65] or channel dimensions [67], as well as hybrid models that capture joint spatial-channel dependencies [44, 59]. Further refinements make use of checkerboard-based decoding [39, 24], codebooks [112], distribution [25, 105, 106], and (lattice) vector quantization [108, 22, 48, 98].
Another body of work seeks to reduce training and inference cost while maintaining compression quality. Notable examples include slimmable sub-networks [88], variable-bit-rate codecs [33, 45, 77], and knowledge distillation strategies [24]. Lightweight decoding has been pursued through shallow or linear decoders [101], while new loss formulations such as causal context [35] or latent decorrelation penalties [1] have also been proposed. Recent studies further introduce rate-distortion-complexity analysis [66, 29, 32, 36], explicitly incorporating computational cost into the optimization objective.
@@ II-B Training Dynamics Approaches for LIC
Although rate-distortion training is often cast as minimizing a scalarized loss $R+\lambda D$, it is fundamentally a multi-objective problem: decreasing rate typically worsens distortion, and vice versa. This observation motivates the adoption of multi-objective optimization (MOO) techniques, which are designed to handle multiple conflicting criteria. One influential MOO approach is the Multiple Gradient Descent Algorithm (MGDA) [15, 85]. MGDA determines an update direction by combining gradients from different objectives with non-negative weights that minimize the squared norm of their sum, subject to a simplex constraint. The resulting direction guarantees improvement for all objectives simultaneously. MOO methods have seen wide adoption in multi-task learning [56, 94, 41], where they are used to balance competing gradients across tasks and mitigate conflicts during training.
@@
Our Perspective. Our work aligns with this emerging line of training-dynamics-based methods and is closely related to Balanced-RD, CMD-LIC, and AuxT. Balanced-RD explicitly regulates the interaction between rate and distortion gradients to promote stable convergence in the R-D trade-off. CMD-LIC accelerates optimization by reducing training space dimensions. AuxT introduces constraints on energy compaction and feature decorrelation to improve convergence behavior.
@@ V-D Empirical Validation
These empirical findings substantiate our central claim: SOAP’s performance gains arise from resolving both intra- and inter-step gradient conflicts in rate-distortion optimization.
@@ VI-C Empirical Validation
These empirical findings support the theoretical perspective in Sec. VI: By coupling channels via Newton-like scaling and rotations, SOAP redistributes variance across directions rather than concentrating it in a few, preventing outlier formation. This yields more regular feature statistics, improving entropy modeling and stabilizing activations, with the downstream benefit of enhanced PTQ robustness. Thus, SOAP not only accelerates training and improves R-D trade-offs but also produces models that are substantially easier to deploy on constrained hardware.
@@ A-D Limitations of Adam for Gradient Alignment
Adam is powerful and widely used, but its effectiveness is limited by its diagonal preconditioner. Because it scales coordinates independently, it cannot exploit off-diagonal curvature that encodes interactions among parameters—precisely what is needed to resolve non-axis-aligned gradient conflicts in multi-objective settings such as rate-distortion (R-D) optimization. The following proposition formalizes a standard local approximation behind this limitation, following [69, 62].
@@ A-E Adam’s Gradient Conflict in a Simplified R-D Setting
Note (positive per-coordinate scalings). The orthogonality conclusion in Proposition 3 is unchanged if Adam’s elementwise normalization introduces arbitrary positive scalings that are independent of the signs of the initialized weights. Concretely, let
@@ A-F Adam’s Behaviour Near a Nondegenerate Optimum
    Frozen second moments. Adam’s second-moment accumulator and damping are (locally) stationary, yielding a fixed positive diagonal matrix $D\succ 0$ and a scalar $c>0$ (absorbing stepsize, bias correction, $\varepsilon$). Hence the Adam update is the diagonally preconditioned step [46, 80, 103]
@@
Since $B$ is symmetric positive definite, inter-step cosines in the $q$-space admit closed forms; cosines in the original coordinates are equivalent up to constants depending only on $\kappa(D)$ (the condition number of D) [76].
@@ A-F1 Inter-step cosine for Adam
the exact inter-step cosine in the $q$-space is
@@
  • Oscillation threshold. If $\eta\,\mu_{1}(u_{t})>1$, the numerator in equation 58 is negative and $\mathcal{S}_{q}<0$. Hence whenever $\eta\,\lambda_{\max}(B)>1$, there exist directions with negative inter-step cosine (flip-flop behaviour).
@@ A-F2 Intra-step cosine for Adam near the optimum
Remarks Nonzero momentum ($\beta_{1}>0$) produces a linear two-term recurrence in the $q$-space; all qualitative conclusions above persist with $B$ replaced by an $O(1)$ affine function of $B$. If $D$ evolves slowly rather than remaining fixed, equation 58-equation 60 apply between preconditioner refreshes with the current $D_{t}$.